Two planes leave an airport. One flying north at 480 mph and the other flying south at 520 mph. How long will it take for the planes to be 3000 miles apart? I know that I have to use the distance formula I just don't know how to set up the problem please help!
Answer by mananth(16946)

North bound 480 mph
south bound 520 mph
..
opposite directions
so add up speed
520+480=1000 mph
distance = 3000 miles
t= d/r
t=3000/1000
t = 3 hours